# The Susceptibility Paradox in Online Social Influence

### v1

@@ Results
To assess whether a more advanced model equipped with a broader range of features can enhance predictive accuracy, we employ an interpretable ensembling machine learning algorithm. This approach also aims to enhance our understanding of the factors that impact users’ susceptibility to social influence. Specifically, we utilize a random forest regression model that incorporates not only susceptibility metrics but also network properties and user metadata. This features set includes friends count, followers count, favorites count (the total number of tweets a user has liked), and statuses count (the total number of tweets a user has posted), along with network properties from the Interaction network such as degree centrality, eigenvector centrality, and clustering coefficient.
We divide users into training and testing sets and use the variance inflation factor (VIF) to assess multicollinearity among the features. In both the Election and Covid datasets, all VIF values are below 5, indicating no significant multicollinearity issues (James et al. 2013). We optimize the random forest regressor using a randomized search on the hyperparameter space to efficiently predict user IAR and SAR scores. We configure the randomized search to sample 100 different parameter settings using 5-fold cross-validation. After identifying the best parameters (see the Appendix for more details), we evaluate each model on a test set using $R^{2}$.
Table 4 provides an overview of the results. The random forest regressor shows promising performance, particularly for the IAR metric. It slightly outperforms the linear regressor, as indicated by the $R^{2}$ difference labeled as $\Delta R^{2}$ in Table 4. This difference is more pronounced for the SAR metric, while it remains quite small (less than 0.05) for the IAR metric, suggesting that IAR can be effectively predicted using solely friends’ IAR, as previously demonstrated in the linear regression model. A study on feature importance further supports this intuition.
Table 4: Summary of the performance of the random forest and linear regressors. $R^{2}$ denotes the coefficient of determination.
@@
Figure 4: Top 5 random forest feature importances in the Covia dataset for IAR and SAR prediction.
@@
Figure 5: SHAP value distributions for SAR predictions in the Covid dataset. Each distribution illustrates the direction and magnitude of the impact of individual features on the model’s output, with features ordered by importance from top to bottom.
@@
To further elucidate the results from feature importance and quantify the specific contribution of each feature to individual predictions, we compute the SHAP (SHapley Additive exPlanations) values (Lundberg and Lee 2017) of the dependent variables predicted in the test set. SHAP values provide a nuanced measure of the impact of each feature on individual instances. Each feature is assigned a SHAP value per instance, which can be aggregated to offer a comprehensive view of the direction and magnitude of its impact on the model’s predictions. Figure 8 shows a SHAP summary plot for the SAR metric in the Covid dataset (additional results are detailed in the Appendix). On the y-axis, features are listed in order of impact, while the x-axis displays the SHAP values, indicating the change in log odds caused by each feature. Points represent individual data instances, colored red (resp. blue) for high (resp. low) feature values.
The most impactful features affecting the SAR predictions in the Covid dataset are friends’ SAR, user statuses count, and user clustering coefficient. The features friends’ SAR and user statuses count have significant opposing impacts on the predictions, indicating that these features can notably increase or decrease the SAR value. Specifically, high values of friends’ SAR tend to increase the SAR score, while higher values of user statuses count tend to decrease it. This relationship between user statuses count and SAR suggests that a high volume of shared tweets correlates with a small rate of spontaneous adoptions. Interestingly, the user clustering coefficient also emerges as a significant factor, where higher clustering coefficients typically decrease SAR, possibly indicating that users with denser social networks may experience fewer spontaneous adoptions. This finding is consistent in the Election dataset (see the Appendix), where user statuses count, friends’ SAR, and user favorites count are the most impactful features.
For IAR prediction (see the Appendix), the most impactful feature is friends’ IAR, with high values of friends’ IAR significantly increasing the IAR score, supporting our previous results related to homophily in IAR.
@@
Users’ level of influence-driven adoption can be effectively predicted by solely leveraging friends’ IAR. In contrast, the predictability of a user’s SAR is significantly enhanced by incorporating additional features, such as counts of tweets and likes. Explainable models reveal that SAR is affected by a wider array of factors, suggesting a more complex interplay of user engagement and preferences in the spontaneous adoption of content.
@@
This study, based on the analysis of millions of tweets, explores the networked dynamics of user susceptibility to online influence from the dual perspective of influence-driven and spontaneous adoption. We present four key insights from our research.
@@
We find that susceptibility is homophilous in social networks, as evidenced by the positive correlations between users’ adoption rates (either influence-driven or spontaneous) and those of their friends. This indicates that individuals prone to influence often connect with similarly susceptible peers. Spontaneous adoption shows significant but lower homophily.
@@
We observe the Generalized Friendship Paradox (GFP) for user susceptibility to influence, termed the Susceptibility Paradox, across two dimensions: First, individuals in a social network tend to be connected to others who are more likely to adopt content due to social influence, as reflected in higher IAR; second, they engage in less spontaneous content adoption, as indicated by lower SAR. Collectively, these observations suggest that “your friends are generally more susceptible to social influence than you.” Additionally, we observe that IAR and SAR are negatively correlated, suggesting that users with higher influence-driven adoption are less likely to share content spontaneously.
@@
While we observe that the GFP holds for IAR but not for SAR, both IAR and SAR demonstrate homophily. These homophilous patterns indicate that individuals are prone to connect with others who exhibit similar rates of influence-driven and spontaneous adoption. This presents two countervailing forces where, although one’s friends typically have higher IAR but lower SAR than oneself, they tend to remain similar in terms of overall susceptibility scores. Comparing the individual- and network-level Generalized Friendship Paradox in homophilous real-world networks versus non-homophilous baseline networks, our findings suggest that the GFP is less pronounced in homophilous networks, where high-IAR (or SAR) nodes tend to connect with other high-IAR (or SAR) nodes. This corroborates Lerman, Yan, and Wu (2016)’s finding that the “majority illusion” is more pronounced in heterogeneous, disassortative networks.
@@
Our analysis indicates that IAR can be accurately predicted solely based on friends’ IAR, whereas SAR is influenced by a broader range of factors. This distinction adds depth to our understanding of susceptibility to influence. On the one hand, our results suggest that influence-driven behaviors (IAR) tend to follow more predictable patterns. This likely stems from the strong link between influence-driven adoption and observable network structures, such as homophily (RQ1) and GFP (RQ2). For instance, studies have shown that targeted marketing and influence campaigns can be effectively strategized based on previous adoption patterns of a consumer’s network (Hill, Provost, and Volinsky 2006). On the other hand, spontaneous behaviors (SAR) are influenced by a myriad of internal and external, often stochastic factors, making them less predictable.
@@
Our research offers practical insights for social media platforms, policymakers, and marketers. The observed homophily in susceptibility suggests that targeted interventions in fields like public health and marketing can effectively target influenceable audiences. By identifying these cohesive, vulnerable groups, intervention strategies can be tailored to mitigate the spread of harmful content and encourage positive behaviors more efficiently (Centola 2011). Further, the observation that the Generalized Friendship Paradox applies to influence-driven adoption could inspire efficient sampling algorithms to identify highly susceptible populations (Alipourfard et al. 2020). Given that IAR is more predictable and that individuals with high IAR are more influenced by their social connections, platforms can specifically target vulnerable audiences by nudging them with accurate, fact-checked content to prevent the endorsement and spread of misinformation.
@@
We acknowledge several limitations in our research. First, our study draws on datasets specifically related to the 2020 US election and the COVID-19 pandemic, capturing only a subset of users over an extended but limited period. Consequently, generalizing our results to broader user networks, different contexts, or other social media platforms necessitates further validation. Second, we infer exposure based on observable user interactions, which might not comprehensively capture exposures within Twitter and across different media. Nonetheless, this method for assessing exposure was tested and validated in prior research (Ferrara and Yang 2015; Rao, Morstatter, and Lerman 2023; Ye et al. 2024). Third, although friends’ IAR explains 36% and 28% of the variation of user IAR in Election and Covid datasets, respectively, we recognize that social influence is a complex phenomenon and may be influenced by various unobserved personal, societal, and cultural factors. Despite these limitations, the consistency of our results across multiple intertwined analyses on distinct datasets supports the reliability and robustness of our findings.
Ethical Statement.
Throughout our research process, we have adhered to stringent ethical standards. We have ensured that no personally identifiable information was utilized, and all analyses were conducted on aggregated data. Additionally, we have thoroughly considered the societal impacts of our research. We acknowledge the potential misuse of our findings by malicious entities to target vulnerable individuals susceptible to manipulation. However, we firmly believe that the significance of our work outweighs these risks. Our research has the potential to inform the development of online systems that prioritize safety and responsibility for all individuals, thereby contributing positively to the integrity of digital environments.
@@
Figure 8: SHAP value distributions for SAR predictions in the Election dataset. The plot illustrates the direction and magnitude of impact of individual features on the model’s output, with features ordered by importance from top to bottom.
Figure 8 shows the SHAP value distributions for SAR predictions in the Election dataset. Figure 9 show the SHAP value distributions for IAR predictions in the Election dataset.
Figure 9: SHAP value distributions for IAR predictions in the Election dataset. The plot illustrates the direction and magnitude of impact of individual features on the model’s output, with features ordered by importance from top to bottom.

### v3

@@ Results
To assess whether a more advanced model equipped with a broader range of features can enhance predictive accuracy, we employ an interpretable ensembling machine learning algorithm. This approach also aims to enhance our understanding of the factors that impact users’ susceptibility to social influence. Specifically, we utilize a random forest regression model that incorporates not only susceptibility metrics but also network properties and user metadata. This feature set includes friends count, followers count, favorites count (the total number of tweets a user has liked), and statuses count (the total number of tweets a user has posted), along with network properties from the Interaction network such as degree centrality, eigenvector centrality, and clustering coefficient.
We divide users into training and testing sets and use the variance inflation factor (VIF) to assess multicollinearity among the features. In both the Election and Covid datasets, all VIF values are below 5, indicating no significant multicollinearity issues (James et al. 2013). We optimize the random forest regressor using a randomized search over the hyperparameter space to efficiently predict user IAR and SAR scores. The randomized search is configured to sample 100 different parameter settings using 5-fold cross-validation. After selecting the optimal parameters (see the Appendix for details), we assess each model’s performance on the test set using the $R^{2}$ metric.
Table 4 provides an overview of the results. The random forest regressor shows promising performance, particularly for the IAR metric. It slightly outperforms the linear regressor, as indicated by the $R^{2}$ difference labeled as $\Delta R^{2}$ in Table 4. This difference is more pronounced for the SAR metric, while it remains quite small (less than 0.05) for the IAR metric, suggesting that IAR can be effectively predicted using solely friends’ IAR, as previously demonstrated in the linear regression model. An analysis of feature importance further supports this intuition.
@@
Table 4: Summary of the performance of the random forest and linear regressors. $R^{2}$ denotes the coefficient of determination and $\Delta R^{2}$ represents the $R^{2}$ difference between random forest and linear regression models.
Figure 6: The top 5 most important features from the random forest model in the COVID dataset for predicting IAR and SAR.
@@
To further elucidate the results from the feature importance analysis and quantify the specific contribution of each feature to individual predictions, we compute the SHAP (SHapley Additive exPlanations) values (Lundberg and Lee 2017) of the dependent variables predicted in the test set. SHAP values provide a nuanced measure of the impact of each feature on individual instances. Each feature is assigned a SHAP value per instance, which can be aggregated to offer a comprehensive view of the direction and magnitude of its impact on the model’s predictions. Figure 7 shows a SHAP summary plot for the SAR and IAR metrics of the Covid dataset (additional results are detailed in the Appendix). On the y-axis, features are listed in order of impact, while the x-axis displays the SHAP values, indicating the change in log odds caused by each feature. Points represent individual data instances, colored red (resp. blue) for high (resp. low) feature values.
(a) SHAP value distributions for SAR predictions.
(b) SHAP value distributions for IAR predictions.
Figure 7: SHAP value distributions for SAR (a) and IAR (b) predictions in the Covid dataset. Each distribution illustrates the direction and magnitude of the impact of individual features on the model’s output, with features ordered by importance from top to bottom.
As shown by Figure 7, the most impactful features affecting the SAR predictions in the Covid dataset are friends’ SAR, user statuses count, and user clustering coefficient. The features friends’ SAR and user statuses count have significant opposing impacts on the predictions, indicating that these features can notably increase or decrease the SAR value. Specifically, high values of friends’ SAR tend to increase the SAR score, while higher values of user statuses count tend to decrease it. This relationship between user statuses count and SAR suggests that a high volume of shared tweets correlates with a small rate of spontaneous adoptions. Interestingly, the user clustering coefficient also emerges as a significant factor, where higher clustering coefficients typically decrease SAR, possibly indicating that users with more cohesive social networks may experience fewer spontaneous adoptions. This finding is consistent in the Election dataset (see the Appendix), where user statuses count, friends’ SAR, and user favorites count are the most impactful features.
For IAR prediction (see Figure 7), the most impactful feature is friends’ IAR, with high values of friends’ IAR significantly increasing the IAR score, supporting our previous results related to homophily and GFP in IAR.
@@
Users’ level of influence-driven adoption can be effectively predicted by solely leveraging friends’ IAR. In contrast, the predictability of a user’s SAR is significantly enhanced by incorporating additional features, such as counts of tweets and likes. Explainable models reveal that SAR is affected by a wider array of factors, suggesting a more complex interplay of user engagement and characteristics in the spontaneous adoption of online content.
Sensitivity Analysis
To address potential sampling biases related to user network construction, we conducted a sensitivity analysis by testing different thresholds for identifying target users. Specifically, we applied thresholds of 5 and 20 shared URLs, in addition to the previously used threshold of 10 URLs, to select the set of target users under scrutiny. This approach allowed us to assess whether varying levels of user activity affected our results. The results, detailed in the Appendix, remained consistent across the different thresholds. Both the IAR and SAR metrics exhibit homophily. Additionally, the Generalized Friendship Paradox (GFP) holds for influence-driven behaviors (IAR) but not for spontaneous actions (SAR), further validating our findings and conclusions.
@@
This study, based on the analysis of millions of tweets, explores the networked dynamics of user susceptibility to online influence from the dual perspective of influence-driven and spontaneous adoption. We present four key insights drawn from our research.
@@
We find that susceptibility is homophilous in social networks, as evidenced by the positive correlations between users’ adoption rates (either influence-driven or spontaneous) and those of their friends. This indicates that influenceable individuals often connect with similarly susceptible peers. Spontaneous adoption shows significant but lower homophily.
@@
We observe the Generalized Friendship Paradox (GFP) for user susceptibility to influence, termed the Susceptibility Paradox, across two dimensions: First, individuals in a social network tend to be connected to others who are more likely to adopt content due to prior exposure, as reflected in a higher IAR. Second, users’ friends engage in less spontaneous content adoption with respect to the users themselves, as indicated by their lower SAR. Collectively, these observations suggest that “your friends are generally more susceptible to social influence than you.” Additionally, we observe that IAR and SAR are negatively correlated, suggesting that users with higher influence-driven adoption are less likely to share content spontaneously.
@@
While we observe that the GFP holds for IAR but not for SAR, both IAR and SAR demonstrate homophily. These homophilous patterns indicate that individuals are prone to connect with others who exhibit similar rates of influence-driven and spontaneous adoption. This reveals two countervailing forces: while an individual’s friends generally have a higher IAR but a lower SAR compared to the individual, their overall susceptibility scores tend to remain similar. Comparing the individual- and network-level GFP in homophilous real-world networks against non-homophilous baseline networks, our findings suggest that the GFP is less pronounced in homophilous networks, where high-IAR (or SAR) nodes tend to connect with other high-IAR (or SAR) nodes. This corroborates the findings of Lerman, Yan, and Wu (2016), which indicate that the “majority illusion” is more pronounced in heterogeneous, disassortative networks.
@@
Our analysis indicates that IAR can be predicted solely based on friends’ IAR, whereas SAR is influenced by a broader range of factors. This distinction adds depth to our understanding of susceptibility to influence. On the one hand, our results suggest that influence-driven behaviors (IAR) tend to follow more predictable patterns. This likely stems from the strong relationship between influence-driven adoption and topological properties, such as homophily (RQ1) and GFP (RQ2), observed in this study. Also, research has shown that targeted marketing and influence campaigns can be effectively strategized based on previous adoption patterns of a consumer’s network (Hill, Provost, and Volinsky 2006). On the other hand, spontaneous behaviors (SAR) are influenced by a myriad of internal and external, often stochastic factors, making them less predictable.
@@
Our research offers practical insights for social media platforms, researchers, policymakers, and marketers. The observed homophily in susceptibility suggests that targeted interventions in fields like public health and marketing can effectively target influenceable audiences. By identifying these cohesive, vulnerable groups, intervention strategies can be tailored to mitigate the spread of harmful content and encourage positive behaviors more efficiently (Centola 2011). Further, the observation that the Generalized Friendship Paradox applies to influence-driven adoption could inspire efficient sampling algorithms to identify highly susceptible populations (Alipourfard et al. 2020).
Predicting users’ IAR and SAR could be valuable in two key scenarios. First, for new users joining an existing social network, their IAR and SAR can be approximated based on interactions with more longevous users, leveraging metrics and attributes of their friends. This allows for an early assessment of a new user’s susceptibility to influence. Second, for less active users or those who rarely share URLs—falling below the threshold required to compute their IAR or SAR—we can infer their susceptibility leveraging the susceptibility metrics of their friends. Overall, this enables social media providers to account for the susceptibility of users with limited activity. Since IAR is more predictable, and individuals with high IAR are more influenced by their social connections, platforms can strategically target vulnerable audiences by nudging them with accurate, fact-checked content to help prevent the endorsement and spread of misinformation (Pennycook et al. 2021).
Future research will extend our susceptibility framework across various media and content modalities. Since the framework is based solely on the dynamics between exposure and adoption, it is platform-agnostic, allowing us to analyze susceptibility on different social media networks. This could involve studying not only URLs in general, but also disentangling specific content types, such as videos, images, and memes. Additionally, our research will explore application scenarios where predicting user susceptibility is critical, such as modeling the impact of coordinated influence operations (Luceri et al. 2024) and exposure to AI-generated disinformation (Augenstein et al. 2024) on content adoption. Validating our model in these domains will further strengthen its real-world applicability and relevance.
@@
We acknowledge several limitations in our research. First, our study draws on datasets specifically related to the 2020 US election and the COVID-19 pandemic, capturing only a subset of users over an extended but limited period. Consequently, generalizing our results to broader user networks, different contexts, or other social media platforms necessitates further validation. Second, we infer exposure based on observable user interactions, which might not comprehensively capture exposures within Twitter and across different media. Nonetheless, this method for assessing exposure was tested and validated in prior research (Ferrara and Yang 2015; Rao, Morstatter, and Lerman 2023; Ye et al. 2024). Third, our network construction method excludes less active users (those sharing fewer than 5, 10, or 20 URLs), which may limit the generalizability of the observed homophily and friendship paradox. Additionally, since we use URL sharing as the adoption criterion, there may be users who rarely share URLs, further restricting the size of the target users under scrutiny. However, our sensitivity analysis shows consistent findings (see related section in the Appendix). Varying the target user thresholds (sharing 5, 10, or 20 URLs) did not affect the significance of our results, indicating the robustness of our findings despite potential selection bias. Fourth, although friends’ IAR explains 36% and 28% of the variation of user IAR in Election and Covid datasets, respectively, we recognize that social influence is a complex phenomenon and may be influenced by various unobserved personal, societal, and cultural factors. Despite these limitations, the consistency of our results across multiple intertwined analyses on distinct datasets supports the reliability and robustness of our findings.
Acknowledgments
Work supported in part by DARPA (contract #HR001121C0169) and the NSF (award #2331722).
@@ Acknowledgments
Figure 10 shows the SHAP value distributions for SAR predictions in the Election dataset. Figure 11 shows the SHAP value distributions for IAR predictions in the Election dataset.
Figure 10: SHAP value distributions for SAR predictions in the Election dataset. Each distribution illustrates the direction and magnitude of the impact of individual features on the model’s output, with features ordered by importance from top to bottom.
Figure 11: SHAP value distributions for IAR predictions in the Election dataset. Each distribution illustrates the direction and magnitude of the impact of individual features on the model’s output, with features ordered by importance from top to bottom.
Sensitivity Analysis with Varying Thresholds for Target User Identification
To address potential sampling biases introduced by the network construction method, we experiment with different URL adoption thresholds to identify target users, specifically testing thresholds of 5 and 20 URLs, in addition to the previously used threshold of 10 URLs. Similarly, we construct three distinct friendship networks based on the types of reciprocal interactions between these target users: Interaction, Retweet, and Mention networks. The resulting node and edge counts for the friendship networks in both datasets, based on these new thresholds, are presented in Table 6.  [^1] The results for the 10 URL threshold ($G_{10}(V,E)$)is available in the main paper.
Consistent with previous findings, the results indicate that both susceptibility metrics (IAR and SAR) exhibit homophily within friendship networks, constructed using thresholds of $5$ and $20$ shared URLs (Table 7). Additionally, IAR demonstrates a stronger degree of homophily compared to SAR. Furthermore, as shown in Table 8, the Generalized Friendship Paradox holds for influence-driven adoption (IAR) but not for spontaneous adoption (SAR) under both thresholds, corroborating our earlier conclusions. These results suggest that our main findings are robust across different sample sizes.
Table 6: Number of nodes ($|V|$) and edges ($|E|$) in the user networks $G_{5}(V,E)$, $G_{10}(V,E)$, and $G_{20}(V,E)$, constructed with URL thresholds of $5$ and $20$, respectively.
Election Covid
$G_{5}(V,E)$ $|V|$ $|E|$ $|V|$ $|E|$
Interaction 98,592 242,856 134,409 300,432
Retweet 56,701 106,893 119,456 198,654
Mention 94,702 148,114 115,584 173,022
Election Covid
$G_{10}(V,E)$ $|V|$ $|E|$ $|V|$ $|E|$
Interaction 49,794 117,384 73,960 163,974
Retweet 25,077 49,679 53,161 90,800
Mention 45,289 72,554 56,060 81,969
Election Covid
$G_{20}(V,E)$ $|V|$ $|E|$ $|V|$ $|E|$
Interaction 45,845 111,402 67,702 154,068
Retweet 23,824 48,588 50,192 87,513
Mention 41,536 67,632 50,695 75,227
Table 7: The susceptibility metrics (IAR and SAR) exhibit homophily within the social network, as evidenced by the weighted average correlation (Wgt. Avg. Corr.) between users and their friends. All correlation coefficients are significant ($p<0.001$).
$G_{5}(V,E)$ Election Covid
Wgt. Avg. Corr. IAR SAR IAR SAR
Interaction 0.512 0.344 0.545 0.352
Retweet 0.520 0.371 0.569 0.356
Mention 0.458 0.302 0.346 0.288
$G_{10}(V,E)$ Election Covid
Wgt. Avg. Corr. IAR SAR IAR SAR
Interaction 0.541 0.356 0.564 0.369
Retweet 0.541 0.391 0.577 0.364
Mention 0.483 0.315 0.372 0.299
$G_{20}(V,E)$ Election Covid
Wgt. Avg. Corr. IAR SAR IAR SAR
Interaction 0.579 0.354 0.577 0.365
Retweet 0.586 0.389 0.589 0.367
Mention 0.506 0.324 0.399 0.294
Table 8: Empirical results for the Generalized Friendship Paradox in three friendship networks derived from the Election dataset are presented. For each susceptibility metric $s$ (IAR and SAR), we calculate the Spearman correlation coefficient $\rho_{ks}$ between degree $k$ and $s$ ($p<0.001$ for all $\rho_{ks}$), average paradox holding probability $P$ and the average susceptibility $\langle s\rangle$ of nodes and their friends $\langle s\rangle_{nn}$.
$G_{5}(V,E)$ Metric $s$ $\rho_{ks}$ Individual-level GFP Network-level GFP
$P$ $\langle s\rangle$ $\langle s\rangle_{nn}$
Interaction IAR 0.188 0.527 0.005 $<$ 0.006
SAR --0.174 0.385 0.889 $>$ 0.876
Retweet IAR 0.302 0.493 0.012 $<$ 0.013
SAR --0.296 0.419 0.826 $>$ 0.812
Mention IAR 0.115 0.532 0.005 $<$ 0.006
SAR --0.103 0.316 0.902 $>$ 0.889
$G_{10}(V,E)$ Metric $s$ $\rho_{ks}$ Individual-level GFP Network-level GFP
$P$ $\langle s\rangle$ $\langle s\rangle_{nn}$
Interaction IAR 0.179 0.541 0.007 $<$ 0.008
SAR --0.166 0.409 0.880 $>$ 0.868
Retweet IAR 0.298 0.501 0.013 $<$ 0.014
SAR --0.294 0.425 0.821 $>$ 0.809
Mention IAR 0.111 0.544 0.005 $<$ 0.006
SAR --0.099 0.383 0.896 $>$ 0.880
$G_{20}(V,E)$ Metric $s$ $\rho_{ks}$ Individual-level GFP Network-level GFP
$P$ $\langle s\rangle$ $\langle s\rangle_{nn}$
Interaction IAR 0.170 0.548 0.007 $<$ 0.009
SAR --0.159 0.419 0.874 $>$ 0.863
Retweet IAR 0.295 0.504 0.013 $<$ 0.014
SAR --0.294 0.428 0.817 $>$ 0.805
Mention IAR 0.099 0.552 0.006 $<$ 0.007
SAR --0.089 0.393 0.890 $>$ 0.874